How many liters of a 10% salt solution must be mixed with 15 liters of a 40% salt solution to obtain a 20% salt solution? I have no idea how to set this eqaution up
Answer by ewatrrr(24785)


Hi
Let x represent represent the amount of 10% solution
Question states***
Mixing together to form (15L+x) 20% solution
.10x + .40*15L = .20(15L+x)
Solving for x
.20*15L = .10x
.20*15L/.10 = x = 30L, the amount of the 10% solution
CHECKING our Answer***
3 + 6 = .20*45 = 9